Fund the length of the diagonal of a rectangular billboard whose sides are 5 ft. and 12 ft. Length = a and width = b given by the Plythagorean Theorem C2 =n a2 + b2.
Answer by checkley71(8403)

5^2+12^2=c^2
25+144=c^2
169=c^2
c=sqrt169
c=13 ft. is the length of the diagonal.